Design a simple ohmmeter and explain the conditions of the ohmmeter circuit when the terminal to which the unknown resistance is connected "open" and "close".
1
Expert's answer
2019-11-04T06:31:42-0500

A simple ohmmeter can be made of 2 constant resistors, 1 variable resistor, a voltage source, and a galvanometer, or a voltmeter. One puts the unknown resistance to x, and adjusts the variable resistor 2 so that the current through the galvanometer is 0 or the voltage between the terminals of the Wheatstone bridge. As soon as 0 is obtained, find the unknown resistance:


"R_x=\frac{R_3R_2}{R_1}."


If the terminals where the x is supposed to be are opened, the ammeter's readings are


"I_G=\frac{ER_1}{R_1R_2+R_2R_3+R_1R_3.}"

If "I_G=0," you should record the value of "R_2" and calculate the unknown resistance according to the first equation.

If there is a voltmeter instead of G, its readings would be


"V_G=\frac{ER_1}{R_1+R_2}."

If "V_G=0," you should record the value of "R_2" ​and calculate the unknown resistance.

Any other readings of "I_G, V_G" mean that the Wheatstone's bridge is unbalanced, and the terminals of "R_x" are closed.